OpenStudy (anonymous):

Analyze the graph of the following function as follows: (a) Find the x- and y-intercepts. (b) Determine the end behavior: find the power function that the graph of f resembles for large values of |x|. (c) Find the maximum number of turning points. f(x)=x^2(x^2-2)(x+2)

OpenStudy (anonymous):

im sorry, its degree 5 lol

OpenStudy (anonymous):



OpenStudy (anonymous):

(a) x-intercept is when f(x)=0 x^2(x^2-2)(x+2)=0 x=0, x=-2, x= +sqrt(2) and x=-sqrt(2) (b) y-intercept is when x=0 f(x)=x^2(x^2-2)(x+2) f(x) = 0

OpenStudy (anonymous):

(c) maximum number of turning points =4. The highest degree of x is x^5 so max turning point is 5-1 = 4

OpenStudy (campbell_st):

there are a maximum of 4 stationary points... which can be turning points or horizontal points of inflexion... pehaps they should have asked for stationary points

OpenStudy (anonymous):

well they didnt ask for exact number of turning points, only the maximum possible